Solve for z, where z is a real number
√(4z+12) = z

Answer by user_dude2008(1862)

√(4z+12) = z

4z+12 = z^2


z^2-4z-12=0

(z-6)(z+2)=0

z=6 or z=-2

only answer is z=6